# Superlinear Convergence of an Interior Point Algorithm on Linear Semi-definite Feasibility Problems with Application to Linear Matrix Inequalities

### v3

@@ 1 Introduction
It proves not an easy task to show superlinear convergence of an implementable interior point algorithm that has polynomial iteration complexity on an SDP with minimal assumptions on the problem and no modifications to the algorithm. This is especially so for the HKM (and its dual) or the NT search direction used in these algorithms, although superlinear convergence is quickly established for the AHO search direction [11, 13], when researchers started focusing their attention on interior point algorithms to solve SDPs. In the literature when the HKM (and its dual) or the NT search direction is used in an interior point algorithm, such as [10], in addition to strict complementarity assumption, nondegeneracy assumption at an optimal solution and modifications to the algorithm, such as solving the corrector-step linear system in an iteration repeatedly instead of only once (“narrowing” the central path neighborhood [^3] We note that this idea is used in [19] to show superlinear convergence of an interior point algorithm on a wide class of conic optimization problems. Also, this is related to the sufficient conditions on superlinear convergence behavior of iterates generated by the algorithm as studied in [22, 23].), need to be imposed for superlinear convergence of the interior point algorithm. In [14], without assuming nondegeneracy, the feasible interior point algorithm considered in the paper is shown to have polynomial iteration complexity and superlinear convergence. However, the algorithm is not easy to implement. The idea behind the algorithm considered in [14] to have superlinear convergence when solving an SDP is to force the centrality measure of the $k^{th}$ iterate to converge to zero as $k$ tends to infinity. This can be enforced in practice by for example, repeatedly solving the corrector-step linear system in an iteration, as in [10], so as to “narrow” the neighborhood of the central path in which iterates lie. Hence, either a polynomial-time interior point algorithm has to be modified or the algorithm is hard to implement for guaranteed superlinear convergence of these algorithms. Otherwise, special structure needs to be imposed on the SDP, such as considering linear semi-definite feasibility problems (LSDFPs), for superlinear convergence [26, 27] [^4] In [26, 27], the search direction used in the interior point algorithm considered is the (dual) HKM and the NT direction respectively.. For the latter, we also require additionally that a suitable initial iterate is chosen. However, the required initial iterate to guarantee superlinear convergence given in [26, 27] is not practical. In this paper, we improve on the results in [26, 27] by proposing an initial iterate that can be obtained in practice (by solving a primal-dual SDP pair), and using this initial iterate to an implementable interior point algorithm on the homogeneous feasibility model of the LSDFP, we have superlinear convergence of the algorithm. The novelty in showing this superlinear convergence result is the nontrivial application of what is known in the literature [26, 27] to show the result. Two key steps that are needed before we can use relevant results in the literature [26, 27] are (i) applying the interior point algorithm on the homogeneous feasibility model of the LSDFP, instead of the LSDFP itself; (ii) suitably reformulating the homogeneous feasibility model of the LSDFP as a semi-definite linear complementary problem (SDLCP) before we can apply these results.
@@ 1.1 Notations
The space of symmetric $n\times n$ matrices is denoted by $S^{n}$. The cone of positive semi-definite (resp., positive definite) symmetric matrices is denoted by $S^{n}_{+}$ (resp. $S^{n}_{++}$). The identity matrix is denoted by $I_{n\times n}$, where $n$ stands for the size of the matrix. We omit the subscript when the size of the identity matrix is clear from the context. The matrix $E^{ij}\in\Re^{n_{1}\times n_{2}}$ is defined to have $1/2$ in its $(i,j)$ and $(j,i)$ entry, and zero everywhere else.
@@ 2 A Linear Semi-definite Feasibility Problem and its Homogeneous Feasibility Model
$\begin{array}[]{ll}\max&b^{T}y\\ {\mbox{subject\ to}}&\sum_{i=1}^{m}y_{i}A_{i}+Y=C,\\ &Y\in S^{n}_{+}.\end{array}$ (8)
Here, $y=(y_{1},\ldots,y_{m})^{T}\in\Re^{m}$. We consider the case when (4) and (8) are feasible in this paper.
A linear semi-definite feasibility problem (LSDFP) is a primal-dual SDP pair (4)-(8) with $C=0$ or $b=0$. WLOG, from now onwards, we consider the LSDFP with $C=0$. Also, we assume that $b_{i}\not=0$, for some $i=1,\ldots,m$, to avoid trivial considerations.
@@
where $z_{j}\in\Re,j=1,\ldots l$, are the decision variables, and $B_{j}\in S^{n},j=0,\ldots,l$, are given. Solving LMIs has wide applicability in diverse areas, as discussed for example in [4]. It is easy to see that (9) can be written as the primal SDP (4) with $C=0$, for suitable $A_{i},b_{i},i=1,\ldots,m$, where we wish to find $X\in S^{n}_{+}$ that solves the primal SDP (4) with $C=0$. Hence, by solving the primal SDP (4) with $C=0$, we are able to solve (9) [^5] It is known that implementable interior point algorithms with polynomial iteration complexities can be used to solve LMIs [4, 5], but to the best of our knowledge, it is unknown whether these algorithms can solve (nonstrict) LMIs with proven good local convergence behavior, which we show in this paper. .
We impose the following assumptions on the LSDFP throughout this paper:
@@
    There exists $(y,Y)\in\Re^{m}\times S^{n}_{++}$ that is feasible to the dual SDP (8) with $C=0$.
@@
We can find $(y,Y)\in\Re^{m}\times S^{n}_{++}$ in Assumption 2.1(a) by using a primal-dual path following interior point algorithm to solve the following primal-dual SDP pair [^6] In fact, by applying the primal-dual path following interior point algorithm on the primal-dual SDP pair, we will know whether Assumption 2.1(a) holds by checking the solution the algorithm converges to.:
[…]
From [27] (see also [23, 26]), we have polynomial iteration complexity and superlinear convergence of an infeasible primal-dual path following interior point algorithm on the above primal-dual SDP pair to obtain its solution. Hence, we are able to get a strictly feasible solution to the dual SDP (8) with $C=0$ efficiently and fast.
In this paper, we apply an infeasible interior point algorithm using the (dual) HKM search direction or the NT search direction to solve the homogeneous feasibility model of the given LSDFP. Since the algorithm is an interior point algorithm, the initial iterate $(X_{0},y_{0},Y_{0})$ to the algorithm needs to satisfy $X_{0},Y_{0}\in S^{n}_{++}$. To prove superlinear convergence of the algorithm on the homogeneous feasibility model of the LSDFP, we further require the initial iterate $(X_{0},y_{0},Y_{0})$ to have $(y_{0},Y_{0})$ feasible to the dual SDP (8) with $C=0$. Although this is an additional requirement on the initial iterate, it can be satisfied in practice, as discussed in Remark 2.2. This superlinear convergence result improves on what is known in the literature [26, 27], where the initial iterate, $(X_{0},y_{0},Y_{0})$, is required to satisfy impractical conditions for superlinear convergence.
@@
Furthermore, observe that a solution to the homogeneous system is readily available and is given by $(X,y,Y,\tau,\kappa)=(0,0,0,0,0)$. However, we cannot derive optimal solutions to (4) and (8) from this. If there exists a solution $(X^{\ast},y^{\ast},Y^{\ast},\tau^{\ast},\kappa^{\ast})$ to (11)-(14) such that $\kappa^{\ast}=0$ and $\tau^{\ast}>0$, then $(X^{\ast}/\tau^{\ast},y^{\ast}/\tau^{\ast},Y^{\ast}/\tau^{\ast})$ is an optimal solution to the primal-dual SDP pair (4)-(8) with zero duality gap. Conversely, if $(X^{\ast},y^{\ast},Y^{\ast})$ is an optimal solution to the primal-dual SDP pair (4)-(8) with zero duality gap, then $(X^{\ast},y^{\ast},Y^{\ast},1,0)$ solves the system (11)-(14). Interior point algorithms when applied to the homogeneous feasibility model can be used to find optimal solutions to the primal-dual SDP pair (4)-(8) by finding a solution $(X^{\ast},y^{\ast},Y^{\ast},\tau^{\ast},\kappa^{\ast})$ to (11)-(14) such that $\kappa^{\ast}=0$ and $\tau^{\ast}>0$. Such an interior point algorithm is discussed in [21], which is also given in Section 4 below where it is specialized to the case when $C=0$. These interior point algorithms necessarily have to be of the infeasible type in that the initial iterate and subsequent iterates generated by the algorithm cannot satisfy (11)-(13). This is so because if an iterate $(X_{k},y_{k},Y_{k},\tau_{k},\kappa_{k})\in S^{n}_{++}\times\Re^{m}\times S^{n}_{++}\times\Re_{++}\times\Re_{++}$ for some $k\geq 0$ satisfies (11)-(13), then (15) holds, which is impossible since $X_{k},Y_{k}\in S^{n}_{++}$ and $\tau_{k},\kappa_{k}>0$.
@@
When specialized to the case when $C=0$, (11)-(14) are given by:
[…]
@@ 3 Homogeneous Feasibility Model as a Semi-definite Linear Complementarity Problem
$\mathcal{A}_{1}(X^{1})+\mathcal{B}_{1}(Y^{1})=q,$ (20)
$X^{1}Y^{1}=0,$ (21)
$X^{1},Y^{1}\in S^{n_{1}}_{+},$ (22)
where $\mathcal{A}_{1},\mathcal{B}_{1}:S^{n_{1}}\rightarrow\Re^{\tilde{n}_{1}}$ are linear operators, $q\in\Re^{\tilde{n}_{1}}$, and $\tilde{n}_{1}=n_{1}(n_{1}+1)/2$. The following assumptions are imposed on the SDLCP, which we show in Proposition 3.2 to hold for the SDLCP representation of the homogeneous feasibility model of the LSDFP. This SDLCP representation will be derived later in the section.
Assumption 3.1
  1. (a)
    System (20)-(22) is monotone. That is, $\mathcal{A}_{1}(X^{1})+\mathcal{B}_{1}(Y^{1})=0$ for $X^{1},Y^{1}\in S^{n_{1}}\Rightarrow{\rm{Tr}}(X^{1}Y^{1})\geq 0$.
  2. (b)
    There exists at least one solution to SDLCP (20)-(22).
  3. (c)
    $\{\mathcal{A}_{1}(X^{1})+\mathcal{B}_{1}(Y^{1})\ ;\ X^{1},Y^{1}\in S^{n_{1}}\}=\Re^{\tilde{n}_{1}}$.
Assumption 3.2
There exist $X^{1},Y^{1}\in S^{n_{1}}_{++}$ such that $\mathcal{A}_{1}(X^{1})+\mathcal{B}_{1}(Y^{1})=q$.
Assumptions 3.1(a), (c) and 3.2 are imposed in [12] where the paper studies feasible interior point algorithms on the SDLCP (20)-(22), while [10, 26] assume Assumption 3.1 in the study of infeasible interior point algorithms on the SDLCP (20)-(22). In the study of infeasible interior point algorithms on the SDLCP (20)-(22), it is not necessary to impose Assumption 3.2. A reason is that we do not need a strictly feasible initial iterate to the algorithm. In this paper, we solve an LSDFP by applying an interior point algorithm on the homogeneous feasibility model (16)-(19). This interior point algorithm is necessarily infeasible as discussed near the end of Section 2. When we express the homogeneous feasibility model as an SDLCP (to be discussed next), the interior point algorithm that is applied to the homogeneous feasibility model can be equivalently applied to the corresponding SDLCP, and by necessity, the algorithm on the SDLCP is infeasible just like that for the homogeneous feasibility model. Therefore, having Assumption 3.2 imposed on the SDLCP is not suitable and in fact can never hold in our case as there cannot exist an $(X,Y)\in S^{n}_{++}\times S^{n}_{++}$ such that (20) holds for the SDLCP obtained from the homogeneous feasibility model. We therefore have Assumption 3.1(b) in its place. In fact, we will see later that when the LSDFP satisfies Assumption 2.1, the SDLCP representation of the homogeneous feasibility model (16)-(19) satisfies Assumption 3.1.
@@
Let $[d_{1}\ {\rm{svec}}(B_{1})^{T}]^{T}$, $d_{1}\not=0$, be orthogonal to the set
$\left\{\left[\begin{array}[]{c}-b_{1}\\ {\mbox{svec}}(A_{1})\end{array}\right],\ldots,\left[\begin{array}[]{c}-b_{m}\\ {\mbox{svec}}(A_{m})\end{array}\right]\right\}.$
Since $d_{1}\not=0$ and $b_{i}\not=0$ for some $i=1,\ldots,m$, $B_{1}$ is necessarily nonzero.
@@
[…]
@@
[…]
@@
[…]
@@
Note that for $(y,Y)$ feasible to the dual SDP (8) with $C=0$, we have the entries on the left-hand side of (58) are equal to zero, except possibly the first entry. This follows from the way $d$ is constructed, and that for such $(y,Y)$, the left-hand side of (33) (or (42)) has all its entries equal to zero, except possibly the first entry. By choosing $B_{1}$ appropriately [^7] Details are provided in Appendix A., we can then apply a result in the literature, namely, Theorem 5.1 in [26], to show superlinear convergence of the interior point algorithm considered in this paper on the homogeneous feasibility model of the LSDFP. This is possible when the initial iterate to the algorithm is from a strictly feasible solution to the dual SDP (8) with $C=0$.
@@
$\left[\begin{array}[]{cccc}\mathcal{A}&-b&0&0\\ 0&0&d&\mathcal{B}\end{array}\right]\left[\begin{array}[]{c}{\mbox{svec}}(X)\\ \tau\\ \kappa\\ {\mbox{svec}}(Y)\end{array}\right]=0.$ (66)
@@
is another way the homogeneous feasibility model of the given LSDFP can be represented.
By switching the position of $d$ and $\mathcal{B}$ in (66), we can rewrite it as
It is easy to convince ourselves that (74), (67) is also a representation of the homogeneous feasibility model of the LSDFP, just like (16)-(19), and (66), (67).
We are now ready to express the homogeneous feasibility model of the LSDFP as SDLCP (20)-(22) by letting $n_{1}=n+1,q=0$, and $\mathcal{A}_{1},\mathcal{B}_{1}$ such that
$\begin{array}[]{lll}(\mathcal{A}_{1}(X^{1}))_{i}&:=&{\rm{Tr}}\left(\left[\begin{array}[]{cc}A_{i}&0\\ 0&-b_{i}\end{array}\right]X^{1}\right),\ \ i=1,\ldots,m,\\ (\mathcal{A}_{1}(X^{1}))_{i}&:=&{\rm{Tr}}(E^{i-m,n+1}X^{1}),\ \ i=m+1,\ldots,m+n,\\ (\mathcal{A}_{1}(X^{1}))_{i}&:=&0,\ \ i=m+n+1,\ldots,\tilde{n}_{1},\end{array}$ (80)
@@
$\begin{array}[]{lll}(\mathcal{B}_{1}(Y^{1}))_{j}&:=&0,\ \ j=1,\ldots,m+n,\\ (\mathcal{B}_{1}(Y^{1}))_{j}&:=&{\rm{Tr}}\left(\left[\begin{array}[]{cc}B_{j-(m+n)}&0\\ 0&d_{j-(m+n)}\end{array}\right]Y^{1}\right),\ \ j=m+n+1,\ldots,\tilde{n}_{1},\end{array}$ (85)
for $X^{1},Y^{1}\in S^{n_{1}}$.

### v4

@@ 1 Introduction
It proves not an easy task to show superlinear convergence of an implementable interior point algorithm that has polynomial iteration complexity on an SDP with minimal assumptions on the problem and no modifications to the algorithm. This is especially so for the HKM (and its dual) or the NT search direction used in these algorithms, although superlinear convergence is quickly established for the AHO search direction [11, 13], when researchers started focusing their attention on interior point algorithms to solve SDPs. In the literature when the HKM (and its dual) or the NT search direction is used in an interior point algorithm, such as [10], in addition to strict complementarity assumption, nondegeneracy assumption at an optimal solution and modifications to the algorithm, such as solving the corrector-step linear system in an iteration repeatedly instead of only once (“narrowing” the central path neighborhood [^3] We note that this idea is used in [19] to show superlinear convergence of an interior point algorithm on a wide class of conic optimization problems. Also, this is related to the sufficient conditions on superlinear convergence behavior of iterates generated by the algorithm as studied in [22, 23].), need to be imposed for superlinear convergence of the interior point algorithm. In [14], without assuming nondegeneracy, the feasible interior point algorithm considered in the paper is shown to have polynomial iteration complexity and superlinear convergence. However, the algorithm is not easy to implement. The idea behind the algorithm considered in [14] to have superlinear convergence when solving an SDP is to force the centrality measure of the $k^{th}$ iterate to converge to zero as $k$ tends to infinity. This can be enforced in practice by for example, repeatedly solving the corrector-step linear system in an iteration, as in [10], so as to “narrow” the neighborhood of the central path in which iterates lie. Hence, either a polynomial-time interior point algorithm has to be modified or the algorithm is hard to implement for guaranteed superlinear convergence of these algorithms. Otherwise, special structure needs to be imposed on the SDP, such as considering linear semi-definite feasibility problems (LSDFPs), for superlinear convergence [26, 27] [^4] In [26, 27], the search direction used in the interior point algorithm considered is the (dual) HKM and the NT direction respectively.. For the latter, we also require additionally that a suitable initial iterate be chosen. However, the required initial iterate to guarantee superlinear convergence given in [26, 27] is not practical, since the conditions required in these papers are not easy to achieve practically. In this paper, we improve on the results in [26, 27] by proposing an initial iterate that can be obtained in practice (by solving a primal-dual SDP pair), and using this initial iterate to an implementable interior point algorithm on the homogeneous feasibility model of the LSDFP, we have superlinear convergence of the algorithm. The novelty in showing this superlinear convergence result is the nontrivial application of what is known in the literature [26, 27] to show the result. Two key steps that are needed before we can use relevant results in the literature [26, 27] are (i) applying the interior point algorithm on the homogeneous feasibility model of the LSDFP, instead of the LSDFP itself; (ii) suitably reformulating the homogeneous feasibility model of the LSDFP as a semi-definite linear complementary problem (SDLCP).
@@ 1.1 Notations
The space of symmetric $n\times n$ matrices is denoted by $S^{n}$. The cone of positive semi-definite (resp., positive definite) symmetric matrices is denoted by $S^{n}_{+}$ (resp. $S^{n}_{++}$). The identity matrix is denoted by $I_{n\times n}$, where $n$ stands for the size of the matrix. We omit the subscript when the size of the identity matrix is clear from the context. The matrix $E^{i,j}\in\Re^{n_{1}\times n_{1}}$ is defined to have $1/2$ in its $(i,j)$ and $(j,i)$ entries, and zero everywhere else.
@@
Given a linear map $\mathcal{A}:S^{n}\rightarrow\Re^{\tilde{n}}$, where $\tilde{n}=n(n+1)/2$, we can view $\mathcal{A}$ as an $\tilde{n}\times\tilde{n}$ matrix, where its $i^{th}$ row is given by ${\rm{svec}}(\mathcal{A}_{i})^{T},i=1,\ldots,\tilde{n}$. Here, $\mathcal{A}_{i}$ is a symmetric matrix of size $n$ which can be easily determined from the given map $\mathcal{A}$.
@@ 2 A Linear Semi-definite Feasibility Problem and its Homogeneous Feasibility Model
$\begin{array}[]{ll}\max&b^{T}y\\ {\mbox{subject\ to}}&\sum_{i=1}^{m}y_{i}A_{i}+Y=C,\\ &Y\in S^{n}_{+}\end{array}$ (8)
where $y=(y_{1},\ldots,y_{m})^{T}\in\Re^{m}$. We consider the case when (4) and (8) are feasible in this paper.
A linear semi-definite feasibility problem (LSDFP) is a primal-dual SDP pair (4)-(8) with $C=0$ or $b=0$. We call the LSDFP with $C=0$, the sLSDFP in this paper. Also, we assume that $b_{i}\not=0$, for some $i=1,\ldots,m$, in the sLSDFP to avoid trivial considerations.
@@
where $z_{j}\in\Re,j=1,\ldots l$, are the decision variables, and $B_{j}\in S^{n},j=0,\ldots,l$, are given. Solving LMIs has wide applicability in diverse areas, as discussed for example in [4]. It is easy to see that (9) can be written as the primal SDP (4) with $C=0$, for suitable $A_{i},b_{i},i=1,\ldots,m$, where we wish to find $X\in S^{n}_{+}$ that solves this problem. Hence, by solving the primal SDP (4) with $C=0$, we are able to solve (9) [^5] It is known that implementable interior point algorithms with polynomial iteration complexities can be used to solve LMIs [4, 5], but to the best of our knowledge, it is unknown whether these algorithms can solve (nonstrict) LMIs with proven good local convergence behavior, which we show in this paper. .
We impose the following assumptions on the sLSDFP throughout this paper:
@@
    There exists $(y,Y)\in\Re^{m}\times S^{n}_{++}$ that is feasible to the dual SDP (8) with $C=0$. We call the dual SDP (8) with $C=0$, the sDSDP from now onwards.
@@
We can find $(y,Y)\in\Re^{m}\times S^{n}_{++}$ in Assumption 2.1(a) by solving the following primal-dual SDP pair with a primal-dual path following interior point algorithm [^6] In fact, by applying the primal-dual path following interior point algorithm on the primal-dual SDP pair, we will know whether Assumption 2.1(a) holds by checking the solution the algorithm converges to.:
$\begin{array}[]{ll}\min&0\\ {\mbox{subject\ to}}&{\mbox{Tr}}(A_{i}X)=0,\ i=1,\ldots,m,\\ &X\in S^{n}_{+},\\ &\\ \max&0\\ {\mbox{subject\ to}}&\sum_{i=1}^{m}y_{i}A_{i}+Y=0,\\ &Y\in S^{n}_{+}.\end{array}$
From [27] (see also [23, 26]), we have polynomial iteration complexity and superlinear convergence of an infeasible primal-dual path following interior point algorithm on the above primal-dual SDP pair to obtain its solution. Hence, we are able to get a strictly feasible solution to the sDSDP efficiently and fast. Extra computational time is certainly needed to generate this stricly feasible solution to the sDSDP. Whether it is worth the computational effort for this when solving practical problems that can be formulated as sLSDFPs depends on the importance to obtain an accurate solution to the given problem fast.
In this paper, we apply an infeasible interior point algorithm using the (dual) HKM search direction or the NT search direction to solve the homogeneous feasibility model of the sLSDFP. Since the algorithm is an interior point algorithm, the initial iterate $(X_{0},y_{0},Y_{0})$ to the algorithm needs to satisfy $X_{0},Y_{0}\in S^{n}_{++}$. To prove superlinear convergence of the algorithm on the homogeneous feasibility model of the sLSDFP, we further require the initial iterate $(X_{0},y_{0},Y_{0})$ to have $(y_{0},Y_{0})$ feasible to the sDSDP. Although this is an additional requirement on the initial iterate, it can be satisfied in practice, as discussed in Remark 2.2. This superlinear convergence result improves on what is known in the literature [26, 27], where the initial iterate, $(X_{0},y_{0},Y_{0})$, is required to satisfy impractical conditions for superlinear convergence.
@@
Furthermore, observe that a solution to the homogeneous system is readily available and is given by $(X,y,Y,\tau,\kappa)=(0,0,0,0,0)$. However, we cannot derive optimal solutions to (4) and (8) from this. If there exists a solution $(X^{\ast},y^{\ast},Y^{\ast},\tau^{\ast},\kappa^{\ast})$ to (11)-(14) such that $\kappa^{\ast}=0$ and $\tau^{\ast}>0$, then $(X^{\ast}/\tau^{\ast},y^{\ast}/\tau^{\ast},Y^{\ast}/\tau^{\ast})$ is an optimal solution to the primal-dual SDP pair (4)-(8) with zero duality gap. Conversely, if $(X^{\ast},y^{\ast},Y^{\ast})$ is an optimal solution to the primal-dual SDP pair (4)-(8) with zero duality gap, then $(X^{\ast},y^{\ast},Y^{\ast},1,0)$ solves the system (11)-(14). Interior point algorithms when applied to the homogeneous feasibility model can be used to find optimal solutions to the primal-dual SDP pair (4)-(8) by finding a solution $(X^{\ast},y^{\ast},Y^{\ast},\tau^{\ast},\kappa^{\ast})$ to (11)-(14) such that $\kappa^{\ast}=0$ and $\tau^{\ast}>0$. Such an interior point algorithm is discussed in [21], which is also given in Section 4 below where it is specialized to the case when $C=0$. These interior point algorithms necessarily have to be of the infeasible type in that the initial iterate and subsequent iterates generated by the algorithm do not always satisfy (11)-(13). This is so because if an iterate $(X_{k},y_{k},Y_{k},\tau_{k},\kappa_{k})\in S^{n}_{++}\times\Re^{m}\times S^{n}_{++}\times\Re_{++}\times\Re_{++}$ for some $k\geq 0$ satisfies (11)-(13), then (15) holds, which is impossible since $X_{k},Y_{k}\in S^{n}_{++}$ and $\tau_{k},\kappa_{k}>0$.
@@
When $C=0$ in (11)-(14), they are given by:
${\mbox{Tr}}(A_{i}X)=b_{i}\tau,\ i=1,\ldots,m,$ (16)
@@ 3 Homogeneous Feasibility Model as a Semi-definite Linear Complementarity Problem
${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}})+{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}})=q,$ (20)
${{\color[rgb]{0,0,0}\hat{X}}}{{\color[rgb]{0,0,0}\hat{Y}}}=0,$ (21)
${{\color[rgb]{0,0,0}\hat{X}}},{{\color[rgb]{0,0,0}\hat{Y}}}\in S^{n_{1}}_{+},$ (22)
where ${{\color[rgb]{0,0,0}\hat{\mathcal{A}}}},{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}:S^{n_{1}}\rightarrow\Re^{\tilde{n}_{1}}$ are linear operators, $q\in\Re^{\tilde{n}_{1}}$, and $\tilde{n}_{1}=n_{1}(n_{1}+1)/2$.
@@
Let $B_{1}\in S^{n}$ and $d_{1}\in\Re,d_{1}\not=0$, such that
$\left[\begin{array}[]{c}d_{1}\\ {\mbox{svec}}(B_{1})\end{array}\right]\perp\left\{\left[\begin{array}[]{c}-b_{1}\\ {\mbox{svec}}(A_{1})\end{array}\right],\ldots,\left[\begin{array}[]{c}-b_{m}\\ {\mbox{svec}}(A_{m})\end{array}\right]\right\}.$
Since $d_{1}\not=0$ and $b_{i}\not=0$ for some $i=1,\ldots,m$, $B_{1}$ is necessarily nonzero [^7] We further require $B_{1}$ to satisfy a more restrictive condition as stated in the appendix..
@@
$\left\{\left[\begin{array}[]{c}-b_{1}\\ {\mbox{svec}}(A_{1})\end{array}\right],\ldots,\left[\begin{array}[]{c}-b_{m}\\ {\mbox{svec}}(A_{m})\end{array}\right],\left[\begin{array}[]{c}d_{1}\\ {\mbox{svec}}(B_{1})\end{array}\right],\left[\begin{array}[]{c}0\\ {\mbox{svec}}(B_{2})\end{array}\right],\ldots,\left[\begin{array}[]{c}0\\ {\mbox{svec}}(B_{\tilde{n}+1-m})\end{array}\right]\right\}$
@@
$\left\{\left[\begin{array}[]{c}-b_{1}\\ {\mbox{svec}}(A_{1})\end{array}\right],\ldots,\left[\begin{array}[]{c}-b_{m}\\ {\mbox{svec}}(A_{m})\end{array}\right]\right\}$
@@
$\left\{\left[\begin{array}[]{c}d_{1}\\ {\mbox{svec}}(B_{1})\end{array}\right],\left[\begin{array}[]{c}0\\ {\mbox{svec}}(B_{2})\end{array}\right],\ldots,\left[\begin{array}[]{c}0\\ {\mbox{svec}}(B_{\tilde{n}+1-m})\end{array}\right]\right\}.$
@@
Note that for $(y,Y)$ feasible to the sDSDP, we have the entries on the left-hand side of (58) are equal to zero, except possibly the first entry. This follows from the way $d$ is constructed, and that for such $(y,Y)$, the left-hand side of (33) (or (42)) has all its entries equal to zero, except possibly the first entry. By choosing $B_{1}$ appropriately [^8] Details are provided in the appendix., we can then apply a result in the literature, namely, Theorem 5.1 in [26], to show superlinear convergence of the interior point algorithm considered in this paper on the homogeneous feasibility model of the sLSDFP. This is possible when the initial iterate to the algorithm is chosen from a strictly feasible solution to the sDSDP.
@@
is another way the homogeneous feasibility model of the sLSDFP can be represented.
It is easy to convince ourselves that (66), (67) is also a representation of the homogeneous feasibility model of the sLSDFP, just like (16)-(19).
We are now ready to express the homogeneous feasibility model of the sLSDFP as SDLCP (20)-(22) by letting $n_{1}=n+1,q=0$, and $^{{\color[rgb]{0,0,0}\hat{\mathcal{A}}}},{{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}$ such that
$\begin{array}[]{lll}({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}}))_{i}&:=&{\rm{Tr}}\left(\left[\begin{array}[]{cc}A_{i}&0\\ 0&-b_{i}\end{array}\right]{{\color[rgb]{0,0,0}\hat{X}}}\right),\ \ i=1,\ldots,m,\\ ({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}}))_{i}&:=&{\rm{Tr}}(E^{i-m,n+1}{{\color[rgb]{0,0,0}\hat{X}}}),\ \ i=m+1,\ldots,m+n,\\ ({{\color[rgb]{0,0,0}\hat{\mathcal{A}}}}({{\color[rgb]{0,0,0}\hat{X}}}))_{i}&:=&0,\ \ i=m+n+1,\ldots,\tilde{n}_{1},\end{array}$ (73)
@@
(^$\begin{array}[]{lll}({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}}))_{j}&:=&0,\ \ j=1,\ldots,m+n,\\ ({{\color[rgb]{0,0,0}\hat{\mathcal{B}}}}({{\color[rgb]{0,0,0}\hat{Y}}}))_{j}&:=&{\rm{Tr}}\left(\left[\begin{array}[]{cc}B_{j-(m+n)}&0\\ 0&d_{j-(m+n)}\end{array}\right]{{\color[rgb]{0,0,0}\hat{Y}}}\right),\ \ j=m+n+1,\ldots,\tilde{n}_{1},\end{array}$ (78)
for ${{\color[rgb]{0,0,0}\hat{X}}},{{\color[rgb]{0,0,0}\hat{Y}}}\in S^{n_{1}}$.
We call this SDLCP, the sSDLCP from now onwards.